OpenStudy (anonymous):

Which of the following points lies on the line whose equation is y = 2x - 3? (0, 3) (1, -1) (-1, 4)

OpenStudy (anonymous):

The easiest way to do this is to check each option one at at time: A) \[y=2x-3\]\[3=2(0)-3\]\[3=0-3\]\[3=-3\] No. B) \[y=2x-3\]\[-1=2(1)-3\]\[-1=2-3\]\[-1=-1\] Yes. C) \[y=2x-3\]\[4=2(-1)-3\]\[4=-2-3\]\[4=-5\] No. The correct answer is B, (1,-1)

OpenStudy (anonymous):

thanks